# EoRA: Fine-tuning-free Compensation for Compressed LLM with Eigenspace Low-Rank Approximation

### v4

@@ Abstract
Abstract: While post-training compression techniques effectively reduce the memory footprint, latency, and power consumption of Large Language Models (LLMs), they often result in noticeable accuracy degradation and remain limited by hardware and kernel constraints that restrict supported compression formats—ultimately reducing flexibility across a wide range of deployment scenarios. In this work, we propose EoRA—a novel, fine-tuning-free method that augments compressed LLMs with low-rank matrices, allowing users to rapidly enhance task-specific performance and freely balance the trade-off between accuracy and computational overhead beyond the constraints of compression formats. EoRA consistently outperforms prior training-free low-rank methods in recovering the accuracy of compressed LLMs, achieving notable accuracy improvements (e.g., $\mathbf{10.84\%}$ on ARC-Challenge, $\mathbf{6.74\%}$ on MathQA, and $\mathbf{6.74\%}$ on GSM8K) for LLaMA3-8B compressed to 3-bit. We also introduce an optimized CUDA kernel, accelerating inference by up to 1.4× and reducing memory overhead through quantizing EoRA. Overall, EoRA offers a prompt solution for improving the accuracy of compressed models under varying user requirements, enabling more efficient and flexible deployment of LLMs. Code is available at https://github.com/NVlabs/EoRA.
@@ 1 Introduction
Figure 1: An overview of our proposed EoRA, which enables swift task-specific accuracy enhancement for compressed LLMs without fine-tuning, using only a small amount of downstream calibration data. At inference time, a single compressed backbone is loaded, while lightweight, task-specific low-rank modules can be dynamically toggled on and off on demand, enabling efficient and flexible deployment. EoRA with rank 128 boosts the accuracy of the LLaMA3-8B model pruned to $2{:}4$ structured sparsity by $4.53\%$, $3.48\%$, and $11.83\%$ on ARC-C, MathQA, and GSM8K, respectively—all achieved within minutes using just 64 calibration samples per task.
Although Large Language Models (LLMs) excel in various tasks, their deployment remains challenging due to high inference costs. Post-training compression methods, like quantization (Frantar et al., 2023; Lin et al., 2024; Liu et al., 2025; Tseng et al., 2024) and pruning (Ma et al., 2023; Frantar and Alistarh, 2023; Sun et al., 2024), aim to reduce computational demands but typically cause accuracy loss or face hardware/kernel constraints, limiting deployment flexibility. For instance, strict hardware-supported formats, such as 2:4 sparsity on NVIDIA GPUs or integer-only quantization kernels, prevent intermediate approaches (e.g., 2.X:4 sparsity or arbitrary-bit quantization) that could offer a more adaptable trade-off between accuracy and latency based on user needs.
To relax these format constraints and improve the accuracy of the compressed models on specified tasks, we formulate a new problem, termed customized compensation: Given a compressed LLM, we attach residual low-rank paths to it to compensate for compression errors and enhance task-specific accuracy, enabling more flexible control over the trade-off between accuracy and compression ratio to accommodate varying user requirements. For example, a user may wish to boost the accuracy of a 2:4 sparsity-pruned model on math reasoning tasks, accepting a modest increase in memory usage and inference latency in return. Importantly, in our problem setting, the weights of the compressed model are not modified during compensation. This enables deployment of a single, general compressed backbone alongside lightweight, task-specific low-rank modules that can be dynamically loaded as needed—allowing for efficient integration with existing multi-adapter inference frameworks (e.g., The vLLM Team ) as illustrated in Figure 1. A naive solution is to apply SVD Li et al. (2024); Yao et al. (2024) for compensation; however, this neglects calibration data and thus fails to enhance task-specific performance. Alternatively, LoRA-based methods, such as Li et al. (2024); Dettmers et al. (2023) require fine-tuning, limiting their applicability for rapid task adaptation. These limitations prompt an important question: “How can we swiftly improve the task-specific accuracy for compressed LLMs without fine-tuning?
@@
We validate the effectiveness of EoRA in boosting the accuracy of compressed LLMs (LLaMA2-7B/13B and LLaMA3-8B) on language generation, commonsense reasoning, and math tasks. Our method consistently outperforms other training-free baselines, especially for aggressively compressed (including pruned, quantized, and both) models (e.g., $\mathbf{2.65\%}$, $\mathbf{3.42\%}$, and $\mathbf{10.99\%}$ improvement on ARC-Challenge, MathQA, and GSM8K when compensating 2:4 pruned LLaMA3-8B). To reduce redundant memory transfer overhead from running low-rank compensation, we design a fused kernel that integrates low-rank and quantization operations, achieving up to 1.4× speedup.
@@
  • Flexible and Task-specific Model Compensation: We propose, fine-tuning-free Eigenspace Low-Rank Approximation (EoRA), a training-free approach that improves the task-specific accuracy of compressed LLMs in minutes using minimal calibration data, while supporting more flexible compression ratios unconstrained by hardware or kernel-imposed format limitations.
@@
2 Preliminaries
Post-training compression aims to compress a well-optimized model by a targeted compression ratio utilizing only a limited set of calibration data. The compression process is often framed as a layer-wise optimization problem, aiming to minimize the layer-wise output difference between the original weight $W_{l}\in\mathbb{R}^{d\times k}$ and the compressed weight $\hat{W}_{l}\in\mathbb{R}^{d\times k}$ for each layer $l$. Then the layer-wise model compression loss can be formed as:
$\operatorname*{arg\,min}_{\hat{W}_{l}}||W_{l}X_{l}-\hat{W}_{l}X_{l}||_{F}$ (1)
@@ 2 Preliminaries
A simple approach to obtain low-rank residual paths that compensate for compression errors is to directly apply Singular Value Decomposition (SVD) (Li et al., 2024; Yao et al., 2024; Li et al., 2025). More specifically, this method relies on a closed-form solution by using SVD to approximate the compression error $\Delta W_{l}=W_{l}-\hat{W}_{l}$ as $\Delta W_{l}\approx U_{l}\Sigma_{l}V_{l}^{T}$, where $\Sigma_{l}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the top-$r$ largest singular value sorted in descending order, and $U_{l}\in\mathbb{R}^{d\times r}$, $V_{l}\in\mathbb{R}^{k\times r}$ are orthonormal matrices, with each column representing the singular vectors corresponding to the singular values in $\Sigma_{l}$. The product of $U_{l}$ and $\Sigma_{l}$ can then be treated as $B_{l}=U_{l}\Sigma_{l}$ with $V_{l}^{T}$ being treated as $A_{l}$. Overall, the error approximation loss can be formulated as:
$\operatorname*{arg\,min}_{B_{l},A_{l}}||\Delta W_{l}-B_{l}A_{l}||_{F}$ (2)
and SVD is applied on $\Delta W_{l}$ to minimize the above equation. However, naively applying SVD to optimize error approximation loss (Eq.2) does not ensure minimization of the layer-wise compression loss (Eq.1) and ignores calibration data, making it ineffective for task-specific accuracy recovery. While LoRA-based methods Li et al. (2024); Dettmers et al. (2023) address this issue, they require fine-tuning and are less suitable for rapid adaptation. This raises a key question: “How can we swiftly improve the task-specific accuracy for compressed LLMs without fine-tuning?”. For simplicity, we omit the subscript $l$, which corresponds to layer $l$ in the following sections.
@@ 3 Method: EoRA
To tackle the challenge of improving task-specific accuracy of compressed LLMs without fine-tuning, we introduce fine-tuning-free Eigenspace Low-Rank Approximation (EoRA)—a method that preserves the efficiency of existing training-free solutions while substantially improving their effectiveness in task-specific accuracy recovery.
@@
First, we propose projecting the compression error into the eigenspace (Stewart, 2001) of the corresponding layer’s input activations, ensuring a direct alignment between the error approximation loss (Eq. 2) and the overall layer-wise model compression loss (Eq. 1). Inspired by the classical Principal Component Analysis (PCA) algorithm, we leverage the eigenvalues of each activation channel as importance scores to indicate the importance of each column after the eigenprojection. This allows us to allocate more low-rank representation capacity to approximate the more critical error elements. Following PCA, we perform the eigendecomposition on $\tilde{X}\tilde{X}^{T}$ where $\tilde{X}\in\mathbb{R}^{k\times n}$ is the average of the input activations over the task-specific calibration set. The eigendecomposition $\tilde{X}\tilde{X}^{T}=Q\Lambda Q^{T}$ is then used to derive the eigenspace projection matrix $Q\in\mathbb{R}^{k\times k}$, whose columns are the eigenvectors, and $\Lambda\in\mathbb{R}^{k\times k}$, which is a diagonal matrix with each diagonal element being the corresponding eigenvalues of the eigenvectors in $Q$. We then propose to project the compression error $\Delta W$ into the eigenspace with the projection matrix $Q^{\prime}=Q\sqrt{\Lambda}$ to obtain the projected error $\Delta W^{\prime}\in\mathbb{R}^{d\times k}=\Delta WQ^{\prime}$. The proposed new error approximation loss, EoRA loss, can be formulated as:
$\operatorname*{arg\,min}_{B^{\prime},A^{\prime}}||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$ (3)
@@
where SVD is applied to approximate $\Delta W\textquoteright$ ’ as $\text{SVD}(\Delta W\textquoteright)\approx U\textquoteright\Sigma\textquoteright V\textquoteright^{T}$, and $\Sigma\textquoteright\in\mathbb{R}^{r\times r}$ contains the top-$r$ singular values. $U\textquoteright\in\mathbb{R}^{d\times r}$ and $V\textquoteright\in\mathbb{R}^{k\times r}$ are orthonormal matrices with columns representing the corresponding singular vectors. Then the low-rank matrices $B^{\prime}$ and $A^{\prime}$ are then assigned as $B^{\prime}=U\textquoteright\Sigma\textquoteright$ ’ and $A^{\prime}=V\textquoteright^{T}$. This loss function ensures that error columns associated with larger eigenvalues are approximated more accurately than those with smaller eigenvalues. We then multiply the low-rank approximation in the eigenspace $\Delta W^{\prime}$ with $Q^{\prime-1}=\sqrt{\Lambda}^{-1}Q^{T}$ to project back to the original space, obtaining the final task-specific compression error approximation as $\Delta W=\Delta W^{\prime}Q^{\prime-1}\approx B^{\prime}A^{\prime}Q^{\prime-1}$. $Q^{\prime}$ is invertible because $Q\textquoteright^{-1}=\sqrt{\Lambda}^{-1}Q^{T}$, and $Q\textquoteright Q\textquoteright^{-1}=Q\sqrt{\Lambda}\sqrt{\Lambda}^{-1}Q^{T}$. Here, the middle term $\sqrt{\Lambda}\sqrt{\Lambda}^{-1}$ simplifies to the identity matrix, and since $Q$ is an orthogonal matrix, $QQ^{T}$ also yields the identity matrix. The product of $A^{\prime}$ and $Q^{\prime-1}$ can be consolidated into a single matrix with the same dimensions as the original $A^{\prime}$, ensuring no additional inference latency as $A=A^{\prime}Q^{\prime-1}$. Then, the forward pass of one linear layer of the compressed model compensated with EoRA for the input activation $X$ can be formulated as:
@@ 7. The final forward pass of current layer becomes $\hat{W}X+B^{\prime}AX$
$\operatorname*{arg\,min}_{B,A}||WX-(\hat{W}+BA)X||_{F}=\operatorname*{arg\,min}_{B,A}||\Delta WX-BAX||_{F}$ (5)
Since the Frobenius norm of a matrix is equal to the square root of its Gram matrix (Sun, 1991; Wang et al., 2025), the minimization problem can be rewritten as:
$&\operatorname*{arg\,min}_{B,A}||\Delta WX-BAX||_{F}=\operatorname*{arg\,min}_{B,A}[\text{trace}((\Delta W-BA)XX^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\end{aligned}$ (6)
@@
$&[\text{trace}((\Delta W-BA)Q\Lambda Q^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\\ &=[\text{trace}((\Delta WQ-BAQ)\Lambda(\Delta WQ-BAQ)^{T})]^{\frac{1}{2}}\end{aligned}$ (7)
@@
Let $Q^{\prime}=Q\sqrt{\Lambda}$, then Eq. 8 becomes:
$&[\text{trace}((\Delta WQ^{\prime}-BAQ^{\prime})(\Delta WQ^{\prime}-BAQ^{\prime})^{T})]^{\frac{1}{2}}\\ &=[\text{trace}((\Delta W^{\prime}-BAQ^{\prime})(\Delta W^{\prime}-BAQ^{\prime})^{T})]^{\frac{1}{2}}\\ &=||\Delta W^{\prime}-BAQ^{\prime}||_{F}\end{aligned}$ (9)
where the square root of the Gram matrix can be transformed back to the corresponding Frobenius norm according to (Sun, 1991). By setting $BAQ^{\prime}=B^{\prime}A^{\prime}$, $||\Delta W^{\prime}-BAQ^{\prime}||_{F}$ becomes $||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$. By the Eckart–Young theorem (Eckart and Young, 1936), the minimization of this Frobenius norm is achieved by running SVD on $\Delta W^{\prime}$, therefore, we prove that minimizing $||\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$ via SVD is equivalent to minimizing Eq. 6, where low-rank approximation of $\Delta W^{\prime}$ is $\text{SVD}(\Delta W^{\prime})=B^{\prime}A^{\prime}$. Note that the above minimization is constrained to the rank of $A^{\prime}$ and $B^{\prime}$.

### v5

@@ Abstract
Abstract: While post-training compression techniques effectively reduce the memory footprint, latency, and power consumption of Large Language Models (LLMs), they often result in noticeable accuracy degradation and remain limited by hardware and kernel constraints that restrict supported compression formats—ultimately reducing flexibility across a wide range of deployment scenarios. In this work, we propose EoRA—a novel, fine-tuning-free method that augments compressed LLMs with low-rank matrices, allowing users to rapidly enhance task-specific performance and freely balance the trade-off between accuracy and computational overhead beyond the constraints of compression formats. EoRA consistently outperforms prior training-free low-rank methods in recovering the accuracy of compressed LLMs, achieving notable accuracy improvements (e.g., $\mathbf{10.84\%}$ on ARC-Challenge, $\mathbf{6.74\%}$ on MathQA, and $\mathbf{11.45\%}$ on GSM8K) for LLaMA3-8B compressed to 3-bit. We also introduce an optimized CUDA kernel, accelerating inference by up to 1.4× and reducing memory overhead through quantizing EoRA. Overall, EoRA offers a prompt solution for improving the accuracy of compressed models under varying user requirements, enabling more efficient and flexible deployment of LLMs. Code is available at https://github.com/NVlabs/EoRA.
Figure 1: An overview of our proposed EoRA, which enables swift task-specific accuracy enhancement for compressed LLMs without fine-tuning, using only a small amount of downstream calibration data. At inference time, a single compressed backbone is loaded, while lightweight, task-specific low-rank modules can be dynamically toggled on and off on demand, enabling efficient and flexible deployment. EoRA with rank 128 boosts the accuracy of the LLaMA3-8B model pruned to 2:42{:}4 structured sparsity by $4.53\%$, $3.48\%$, and $11.83\%$ on ARC-C, MathQA, and GSM8K, respectively—all achieved within minutes using just 64 calibration samples per task.
@@ 1 Introduction
Although Large Language Models (LLMs) excel in various tasks, their deployment remains challenging due to high inference costs. Post-training compression methods, like quantization (frantar2022gptq; lin2024awq; liu2024spinquant; tseng2024quip) and pruning (ma2023llm; frantar2023sparsegpt; sun2023simple), aim to reduce computational demands but typically cause accuracy loss or face hardware/kernel constraints, limiting deployment flexibility. For instance, strict hardware-supported formats, such as 2:4 sparsity on NVIDIA GPUs or integer-only quantization kernels, prevent intermediate approaches (e.g., 2.X:4 sparsity or arbitrary-bit quantization) that could offer a more adaptable trade-off between accuracy and latency based on user needs.
To relax these format constraints and improve the accuracy of the compressed models on specified tasks, we formulate a new problem, termed customized compensation: Given a compressed LLM, we attach residual low-rank paths to it to compensate for compression errors and enhance task-specific accuracy, enabling more flexible control over the trade-off between accuracy and compression ratio to accommodate varying user requirements. For example, a user may wish to boost the accuracy of a 2:4 sparsity-pruned model on math reasoning tasks, accepting a modest increase in memory usage and inference latency in return. Importantly, in our problem setting, the weights of the compressed model are not modified during compensation. This enables deployment of a single, general compressed backbone alongside lightweight, task-specific low-rank modules that can be dynamically loaded as needed—allowing for efficient integration with existing multi-adapter inference frameworks (e.g., vLLM (multilora)) as illustrated in Figure 1. A naive solution is to apply SVD (liloftq; yao2024zeroquantv2) for compensation; however, this neglects calibration data and thus fails to enhance task-specific performance. Alternatively, LoRA-based methods, such as liloftq; dettmers2024qlora, require fine-tuning, limiting their applicability for rapid task adaptation. These limitations prompt an important question: “How can we swiftly improve the task-specific accuracy for compressed LLMs without fine-tuning?
@@
We validate the effectiveness of EoRA in boosting the accuracy of compressed LLMs (LLaMA2-7B/13B and LLaMA3-8B) on language generation, commonsense reasoning, and math tasks. Our method consistently outperforms other fine-tuning-free baselines, especially for aggressively compressed (including pruned, quantized, and both) models (e.g., $\mathbf{2.65\%}$, $\mathbf{3.42\%}$, and $\mathbf{10.99\%}$ improvement over ZeroQuant-V2 on ARC-Challenge, MathQA, and GSM8K when compensating 2:4 pruned LLaMA3-8B). To reduce redundant memory transfer overhead from running low-rank compensation, we design a fused kernel that integrates low-rank and quantization operations, achieving up to 1.4× speedup.
@@
  • Flexible and Task-specific Model Compensation: We propose, fine-tuning-free Eigenspace Low-Rank Approximation (EoRA), a fine-tuning-free approach that improves the task-specific accuracy of compressed LLMs in minutes using minimal calibration data, while supporting more flexible compression ratios unconstrained by hardware or kernel-imposed format limitations.
@@
2 Preliminaries: Post-training Compression
Post-training compression aims to compress a well-trained model by a targeted compression ratio, utilizing only a limited set of calibration data. The compression process is often framed as a layer-wise optimization problem, aiming to minimize the layer-wise output difference between the original weight $W_{l}\in\mathbb{R}^{d\times k}$ and the compressed weight $\hat{W}_{l}\in\mathbb{R}^{d\times k}$ for each layer $l$. Then the layer-wise model compression loss can be formed as:
[…]
@@ 2 Preliminaries: Post-training Compression
A simple approach to obtain low-rank residual paths that compensate for compression errors is to directly apply Singular Value Decomposition (SVD) (liloftq; yao2024zeroquantv2; li2024svdqunat). More specifically, this method relies on a closed-form solution by using SVD to approximate the compression error $\Delta W_{l}=W_{l}-\hat{W}_{l}$ as $\Delta W_{l}\approx U_{l}\Sigma_{l}V_{l}^{T}$, where $\Sigma_{l}\in\mathbb{R}^{r\times r}$ is a diagonal matrix containing the top-$r$ largest singular value sorted in descending order, and $U_{l}\in\mathbb{R}^{d\times r}$, $V_{l}\in\mathbb{R}^{k\times r}$ are orthonormal matrices, with each column representing the singular vectors corresponding to the singular values in $\Sigma_{l}$. The product of $U_{l}$ and $\Sigma_{l}$ can then be treated as $B_{l}=U_{l}\Sigma_{l}$ with $V_{l}^{T}$ being treated as $A_{l}$. Overall, the error approximation loss can be formulated as:
[…]
and SVD is applied on $\Delta W_{l}$ to minimize the above equation. However, naively applying SVD to optimize error approximation loss (Eq.2) does not ensure minimization of the layer-wise compression loss (Eq.1) and ignores calibration data, making it ineffective for task-specific accuracy recovery. While LoRA-based methods (liloftq; dettmers2024qlora) address this issue, they require fine-tuning and are less suitable for rapid adaptation. This raises a key question: “How can we swiftly improve the task-specific accuracy for compressed LLMs without fine-tuning?”. For simplicity, we omit the subscript $l$, which corresponds to layer $l$ in the following sections.
@@ 3 Method: EoRA
To tackle the challenge of improving task-specific accuracy of compressed LLMs without fine-tuning, we introduce fine-tuning-free Eigenspace Low-Rank Approximation (EoRA)—a method that preserves the efficiency of existing fine-tuning-free solutions while substantially improving their effectiveness in task-specific accuracy recovery.
@@
First, we propose projecting the compression error into the eigenspace (stewart2001matrix) of the corresponding layer’s input activations, ensuring a direct alignment between the error approximation loss (Eq. 2) and the overall layer-wise model compression loss (Eq. 1). Inspired by the classical Principal Component Analysis (PCA) algorithm, we leverage the eigenvalues of each activation channel as importance scores to indicate the importance of each column after the eigenprojection. This allows us to allocate more low-rank representation capacity to approximate the more critical error elements. Following PCA, we perform the eigendecomposition on $\tilde{X}\tilde{X}^{T}$ where $\tilde{X}\in\mathbb{R}^{k\times n}$ is the average of the input activations over the task-specific calibration set. The eigendecomposition $\tilde{X}\tilde{X}^{T}=Q\Lambda Q^{T}$ is then used to derive the eigenspace projection matrix $Q\in\mathbb{R}^{k\times k}$, whose columns are the eigenvectors, and $\Lambda\in\mathbb{R}^{k\times k}$, which is a diagonal matrix with each diagonal element being the corresponding eigenvalues of the eigenvectors in $Q$. We then propose to project the compression error $\Delta W$ into the eigenspace with the projection matrix $Q^{\prime}=Q\sqrt{\Lambda}$ to obtain the projected error $\Delta W^{\prime}\in\mathbb{R}^{d\times k}=\Delta WQ^{\prime}$. The proposed new error approximation loss, EoRA loss, can be formulated as:
[…]
@@
where SVD is applied to approximate $\Delta W^{\prime}$ as $\text{SVD}(\Delta W^{\prime})\approx U^{\prime}\Sigma^{\prime}V^{\prime T}$, and $\Sigma^{\prime}\in\mathbb{R}^{r\times r}$ contains the top-$r$ singular values. $U^{\prime}\in\mathbb{R}^{d\times r}$ and $V^{\prime}\in\mathbb{R}^{k\times r}$ are orthonormal matrices with columns representing the corresponding singular vectors. Then the low-rank matrices $B^{\prime}$ and $A^{\prime}$ are then assigned as $B^{\prime}=U^{\prime}\Sigma^{\prime}$ and $A^{\prime}=V^{\prime T}$. This loss function ensures that error columns associated with larger eigenvalues are approximated more accurately than those with smaller eigenvalues. We then multiply the low-rank approximation in the eigenspace $\Delta W^{\prime}$ with $Q^{\prime-1}=\sqrt{\Lambda}^{-1}Q^{T}$ to project back to the original space, obtaining the final task-specific compression error approximation as $\Delta W=\Delta W^{\prime}Q^{\prime-1}\approx B^{\prime}A^{\prime}Q^{\prime-1}$. $Q^{\prime}$ is invertible because $Q^{\prime-1}=\sqrt{\Lambda}^{-1}Q^{T}$, and $Q^{\prime}Q^{\prime-1}=Q\sqrt{\Lambda}\sqrt{\Lambda}^{-1}Q^{T}$. Here, the middle term $\sqrt{\Lambda}\sqrt{\Lambda}^{-1}$ simplifies to the identity matrix, and since $Q$ is an orthogonal matrix, $QQ^{T}$ also yields the identity matrix. The product of $A^{\prime}$ and $Q^{\prime-1}$ can be consolidated into a single matrix with the same dimensions as the original $A^{\prime}$, ensuring no additional inference latency as $A=A^{\prime}Q^{\prime-1}$. Then, the forward pass of one linear layer of the compressed model compensated with EoRA for the input activation $X$ can be formulated as:
@@ 7. The final forward pass of current layer becomes $\hat{W}X+B^{\prime}AX$
[…]
Since the Frobenius norm of a matrix is equal to the square root of its Gram matrix (sun1991proof; wang2024svdllm), the minimization problem can be rewritten as:
$\begin{aligned} &\operatorname*{arg\,min}_{B,A}||\Delta WX-BAX||_{F}=\operatorname*{arg\,min}_{B,A}[\text{trace}((\Delta W-BA)XX^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\end{aligned}$ (6)
@@
$\begin{aligned} &[\text{trace}((\Delta W-BA)Q\Lambda Q^{T}(\Delta W-BA)^{T})]^{\frac{1}{2}}\\ &=[\text{trace}((\Delta WQ-BAQ)\Lambda(\Delta WQ-BAQ)^{T})]^{\frac{1}{2}}\end{aligned}$ (7)
@@
$\begin{aligned} [\text{trace}((\Delta WQ\sqrt{\Lambda}-BAQ\sqrt{\Lambda})(\Delta WQ\sqrt{\Lambda}-BAQ\sqrt{\Lambda})^{T})]^{\frac{1}{2}}\end{aligned}$ (8)
Let $Q^{\prime}=Q\sqrt{\Lambda}$, then Eq. 8 becomes:
@@
where the square root of the Gram matrix can be transformed back to the corresponding Frobenius norm according to (sun1991proof). By setting $BAQ^{\prime}=B^{\prime}A^{\prime}$, $\Delta W^{\prime}-BAQ^{\prime}||_{F}$ becomes $\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$. By the Eckart–Young theorem (Eckart_Young_1936), the minimization of this Frobenius norm is achieved by running SVD on $\Delta W^{\prime}$, therefore, we prove that minimizing $\Delta W^{\prime}-B^{\prime}A^{\prime}||_{F}$ via SVD is equivalent to minimizing Eq. 6, where low-rank approximation of $\Delta W^{\prime}$ is $\text{SVD}(\Delta W^{\prime})=B^{\prime}A^{\prime}$. Note that the above minimization is constrained to the rank of $A^{\prime}$ and $B^{\prime}$.